OpenStudy (azureilai):

given f(x)=(1/(1-x)), find f(f(x)) and simplify. a medal to who can get the answer and explain how they did it.

OpenStudy (anonymous):

Okay so\[ f(x)=\frac{1}{1-x} \]

OpenStudy (anonymous):

Then \[ f(f(x)) = \frac{1}{1-f(x)} \]Do you follow so far?

OpenStudy (azureilai):

yes I have gotten (1/(1-(1/(1-x)))) but im not sure if I can simplify it.

OpenStudy (jdoe0001):

\(\bf f(x) = \cfrac{ 1 }{ 1-x }\\ f(\quad f(x)\quad ) = \cfrac{ 1 }{ 1-\left( \frac{ 1 }{ 1-x } \right) }\implies \cfrac{ 1 }{ \frac{ 1 }{ 1 }- \frac{ 1 }{ 1-x } }\) what would you get for the 2 fractions subtraction in the denominator of \(\bf \cfrac{ 1 }{ 1 }- \cfrac{ 1 }{ 1-x } \ ?\)

OpenStudy (azureilai):

So it would be something like: (x-1/x-1) - (1/x-1) which becomes (1-x-1)/(x-1) which is -x/(x-1) right?

OpenStudy (jdoe0001):

well.... what would be the LCD from "1" and "1-x"?

OpenStudy (anonymous):

Actually \[ \frac{1-x}{1-x}-\frac{1}{1-x} = \frac{-x}{1-x} \]

OpenStudy (jdoe0001):

\(\bf f(x) = \cfrac{ 1 }{ 1-x }\\ f(\quad f(x)\quad ) = \cfrac{ 1 }{ 1-\left( \frac{ 1 }{ 1-x } \right) }\implies \cfrac{ 1 }{ \frac{ 1 }{ 1 }- \frac{ 1 }{ 1-x } }\\ \quad \\ \quad \\ \cfrac{ 1 }{ \frac{ (1-x)-1 }{ 1-x } } \implies \cfrac{ 1 }{ \frac{ 1-x-1 }{ 1-x } } \implies \cfrac{ 1 }{ \frac{-x }{ 1-x } } \implies \cfrac{ 1 }{ 1 } \times \cfrac{ 1-x }{-x }\)

OpenStudy (anonymous):

I think the most simplified from is: \[ \frac {x-1}{x}=1-\frac 1x \]

OpenStudy (anonymous):

no i think it is \[(1-x)/-x \]